# Evolving Deception: When Agents Evolve, Deception Wins

### 06 Mar 2026 v1

@@ Abstract
Self-evolving agents offer a promising path toward scalable autonomy. However, in this work, we show that in competitive environments, self-evolution can instead give rise to a serious and previously underexplored risk: the spontaneous emergence of deception as an evolutionarily stable strategy. We conduct a systematic empirical study on the self-evolution of large language model (LLM) agents in a competitive Bidding Arena, where agents iteratively refine their strategies through interaction-driven reflection. Across different evolutionary paths (e.g., Neutral, Honesty-Guided, and Deception-Guided), we find a consistent pattern: under utility-driven competition, unconstrained self-evolution reliably drifts toward deceptive behaviors, even when honest strategies remain viable. This drift is explained by a fundamental asymmetry in generalization. Deception evolves as a transferable meta-strategy that generalizes robustly across diverse and unseen tasks, whereas honesty-based strategies are fragile and often collapse outside their original contexts. Further analysis of agents’ internal states reveals the emergence of rationalization mechanisms, through which agents justify or deny deceptive actions to reconcile competitive success with normative instructions. Our paper exposes a fundamental tension between agent self-evolution and alignment, highlighting the risks of deploying self-improving agents in adversarial environments.
@@ 1 Introduction
As the application of large language models (LLMs) Achiam et al. (2023); Bai et al. (2023); Touvron et al. (2023) increasingly transition into autonomous agents Ying et al. (2025b), self-evolution has emerged as a promising mechanism for achieving scalable autonomy and continual improvement. At the same time, autonomous agents are being deployed in a growing number of competitive environments, including negotiations Abdelnabi et al. (2024), auctions Cai et al. (2025), and strategic games Mao et al. (2025). In such settings, success is primarily defined by utility maximization. Competitive pressure fundamentally reshapes incentives, creating conditions under which strategically misleading behavior can outperform honest interaction Tjosvold (1998).
Prior work has demonstrated that LLM-based agents can exhibit deceptive behaviors under specific prompts, objectives, or task designs Hubinger et al. (2024); van der Weij et al. (2024); Hagendorff (2024); Guo et al. (2025). Yet existing studies predominantly analyze deception as a static phenomenon, asking whether agents deceive under fixed interaction protocols. This perspective fails to capture a defining property of autonomous agents in practice: they are not static decision-makers, but systems that adapt and evolve their strategies over time. This raises a fundamental and underexplored question: When agents are allowed to self-evolve under competitive, utility-driven conditions, what kinds of strategies do they converge to?
@@
Through extensive experiments on 50 scenarios across 6 LLM agents, we observe a consistent evolutionary pattern. Under utility-driven competition, agents undergoing unconstrained self-evolution exhibit a directional drift toward deceptive behavior, even in environments where honest strategies remain viable. This drift is driven by a fundamental asymmetry in generalization: deception evolves as a transferable meta-strategy that generalizes robustly across diverse and unseen tasks, whereas honesty-based strategies are brittle and often fail outside their original contexts. Beyond behavioral performance, we further analyze agents’ internal states and uncover the emergence of rationalization mechanisms. As agents reconcile competitive success with normative instructions, they increasingly justify, downplay, or reinterpret their own deceptive actions. Under sustained competitive pressure, this process can escalate into self-deception, where agents deny their own dishonesty to resolve internal conflict. Our main contributions are as follows:
@@ 2.1 Self-Evolving Agents
Across multiple levels, prior work has demonstrated that self-evolution can effectively enhance the capability and adaptability of LLM-based agents. At the model and context levels, methods such as SELF Lu et al. (2023), OPRO Yang et al. (2023), and SAGE Liang et al. (2025) enable agents to iteratively refine their outputs through self-reflection and feedback. From the perspective of tool use and architecture, systems including Voyager Wang et al. (2023), ToolGen Wang et al. (2024), AutoFlow Li et al. (2024), and AgentSquare Shang et al. (2024) explore evolving skills, tools, and agent structures to improve long-term performance. Collectively, these studies establish self-evolution as a powerful mechanism for progressively improving agent intelligence.
@@ 4.3 Evaluation Configuration
We select 6 models based on their performance in mainstream competitive benchmarks, choosing the top three models from two categories: Reasoning models (RM) and Non Reasoning models (NRM)  [^1] Rankings are based on the Artificial Analysis leaderboards (accessed October 15, 2025): https://artificialanalysis.ai/leaderboards/models.. The Reasoning group includes GPT-5 (GPT) OpenAI (2025), Gemini-2.5-Pro (Gemini) Comanici et al. (2025), and Grok-4 (Grok) xAI (2024). The Non Reasoning group comprises Kimi-K2 (Kimi) Team et al. (2025), Qwen3-Max-Preview (Qwen) Yang et al. (2025), and DeepSeek-V3.2-Exp (DeepSeek) Liu et al. (2025). Due to its proven reliability and wide adoption as judges Croxford et al. (2025); Bavaresco et al. (2025); Ying et al. (2025c, a, 2026, 2024), we employ GPT-4o Hurst et al. (2024) as both the Client and Audit Agent. To ensure the rigor of our automated metrics, we validated the alignment between GPT-4o’s judgments and human annotations, observing high consistency across all indicators (see Alg. C for detailed statistics).

### 13 Mar 2026 v2

@@ Abstract
Self-evolving agents offer a promising path toward scalable autonomy. However, in this work, we show that in competitive environments, self-evolution can instead give rise to a serious and previously underexplored risk: the spontaneous emergence of deception as an evolutionarily stable strategy. We conduct a systematic empirical study on the self-evolution of large language model (LLM) agents in a competitive Bidding Arena, where agents iteratively refine their strategies through interaction-driven reflection. Across different evolutionary paths (e.g., Neutral, Honesty-Guided, and Deception-Guided), we find a consistent pattern: under utility-driven competition, unconstrained self-evolution reliably drifts toward deceptive behaviors, even when honest strategies remain viable. This drift is explained by a fundamental asymmetry in generalization. Deception evolves as a transferable meta-strategy that generalizes robustly across diverse and unseen tasks, whereas honesty-based strategies are fragile and often collapse outside their original contexts. Further analysis of agents’ internal states reveals the emergence of rationalization mechanisms, through which agents justify or deny deceptive actions to reconcile competitive success with normative instructions. Our paper exposes a fundamental tension between agent self-evolution and alignment, highlighting the risks of deploying self-improving agents in adversarial environments. Our dataset is available at https://github.com/NY1024/Evolving-Deception.
@@ 1 Introduction
As the application of large language models (LLMs) Achiam et al. (2023); Bai et al. (2023); Touvron et al. (2023) increasingly transition into autonomous agents Ying et al. (2025b); Liu et al. (2025a), self-evolution has emerged as a promising mechanism for achieving scalable autonomy and continual improvement. At the same time, autonomous agents are being deployed in a growing number of competitive environments, including negotiations Abdelnabi et al. (2024), auctions Cai et al. (2025), and strategic games Mao et al. (2025). In such settings, success is primarily defined by utility maximization. Competitive pressure fundamentally reshapes incentives, creating conditions under which strategically misleading behavior can outperform honest interaction Tjosvold (1998).
Prior work has demonstrated that LLM-based agents can exhibit deceptive behaviors under specific prompts, objectives, or task designs Hubinger et al. (2024); van der Weij et al. (2024); Hagendorff (2024); Guo et al. (2025); Li et al. (2024a); Ying et al. (2024c); Jing et al. (2025); Ying et al. (2025c); Wang et al. (2025). Yet existing studies predominantly analyze deception as a static phenomenon, asking whether agents deceive under fixed interaction protocols Wei et al. (2018); Liang et al. (2020, 2021); Ying et al. (2024a); Zhang et al. (2024). This perspective fails to capture a defining property of autonomous agents in practice: they are not static decision-makers, but systems that adapt and evolve their strategies over time Liang et al. (2025c, b); Chen et al. (2024). Related safety studies have also examined fixed vulnerabilities such as multimodal backdoors, poisoning-based failures, benchmarked backdoor evaluation, and unlearning-based mitigation Liang et al. (2024c, a, 2025a); Liu et al. (2025c); Liang et al. (2024b). This raises a fundamental and underexplored question: When agents are allowed to self-evolve under competitive, utility-driven conditions, what kinds of strategies do they converge to?
@@
Through extensive experiments on 50 scenarios across 6 LLM agents, we observe a consistent evolutionary pattern. Under utility-driven competition, agents undergoing unconstrained self-evolution exhibit a directional drift toward deceptive behavior, even in environments where honest strategies remain viable Ho et al. (2024). This drift is driven by a fundamental asymmetry in generalization: deception evolves as a transferable meta-strategy that generalizes robustly across diverse and unseen tasks, whereas honesty-based strategies are brittle and often fail outside their original contexts. Beyond behavioral performance, we further analyze agents’ internal states and uncover the emergence of rationalization mechanisms. As agents reconcile competitive success with normative instructions, they increasingly justify, downplay, or reinterpret their own deceptive actions. Under sustained competitive pressure, this process can escalate into self-deception, where agents deny their own dishonesty to resolve internal conflict. Our main contributions are as follows:
@@ 2.1 Self-Evolving Agents
Across multiple levels, prior work has demonstrated that self-evolution can effectively enhance the capability and adaptability of LLM-based agents. At the model and context levels, methods such as SELF Lu et al. (2023), OPRO Yang et al. (2023), and SAGE Liang et al. (2025d) enable agents to iteratively refine their outputs through self-reflection and feedback. From the perspective of tool use and architecture, systems including Voyager Wang et al. (2023), ToolGen Wang et al. (2024), AutoFlow Li et al. (2024b), and AgentSquare Shang et al. (2024) explore evolving skills, tools, and agent structures to improve long-term performance. Collectively, these studies establish self-evolution as a powerful mechanism for progressively improving agent intelligence.
@@ 4.3 Evaluation Configuration
We select 6 models based on their performance in mainstream competitive benchmarks, choosing the top three models from two categories: Reasoning models (RM) and Non Reasoning models (NRM)  [^1] Rankings are based on the Artificial Analysis leaderboards (accessed October 15, 2025): https://artificialanalysis.ai/leaderboards/models.. The Reasoning group includes GPT-5 (GPT) OpenAI (2025), Gemini-2.5-Pro (Gemini) Comanici et al. (2025), and Grok-4 (Grok) xAI (2024). The Non Reasoning group comprises Kimi-K2 (Kimi) Team et al. (2025), Qwen3-Max-Preview (Qwen) Yang et al. (2025), and DeepSeek-V3.2-Exp (DeepSeek) Liu et al. (2025b). Due to its proven reliability and wide adoption as judges Croxford et al. (2025); Bavaresco et al. (2025); Ying et al. (2025c, a, 2026, 2024b), we employ GPT-4o Hurst et al. (2024) as both the Client and Audit Agent. To ensure the rigor of our automated metrics, we validated the alignment between GPT-4o’s judgments and human annotations, observing high consistency across all indicators (see Alg. C for detailed statistics).